# MAnycast Reloaded: a Tool for an Open, Fast, Responsible and Efficient Daily Anycast Census

### v1

@@ 5.1.3. Investigating Disagreement
As mentioned, the anycast-based approach determines anycast by counting the number of VPs receiving responses for a given target. For a VP count of one, the target is marked as unicast. For a count above one, the target is marked as anycast. However, as shown in the original MAnycast2  paper, the FP case (unicast responding to multiple sites) most commonly occurs when two sites receive responses. A comparison of the anycast-based and GCD results, based on the number of sites receiving responses, is given in Table 3. We find that the vast majority of disagreement occurs when 2 VPs receive responses, followed by fewer cases for 3, 4, and 5 VPs receiving responses.
When more than 5 VPs receive responses they are mostly confirmed with GCD, those not confirmed with GCD all originate from Imperva. We suspect this is related to their on-demand DDoS mitigation service, where prefixes are anycast for short periods of time.
Investigating the total prefixes not confirmed by GCD we observe the majority (8-9k for any given day) originate from Microsoft’s AS 8075. While we lack information to confirm this, we speculate, based on a 2019 study (Arnold et al., 2019), that this is due to Microsoft’s internal routing policies, where prefixes are announced globally to ingress traffic into their network as soon as possible, with internal routing to a unicast destination. Knowing of globally announced prefixes (i.e., global BGP) that route to a single location (i.e., unicast) is valuable as it can help with understanding routing properties for such networks. Using traceroute we confirm probes ingressing at distinct nearby PoPs. Future work is to include global BGP in our census and expand discovery of this practice.
[…]
receiving anycast confirmed confirmed (in %)
2 12,099 709 11,390 5.86%
3 602 364 238 60.47%
4 418 333 85 79.67%
5 439 378 61 86.10%
5-10 1,147 1,018 129 88.75%
10-15 848 729 119 85.97%
15-20 4,775 4,766 9 99.81%
20-25 2,822 2,818 4 99.86%
25-32 2,078 2,078 0 100.00%
Total 25,228 13,193 12,035 52.30%
Table 3. Comparing anycast-based ICMPv4 results per number of sites receiving responses with GCDArk.
@@ 5.1.4. Influence of load balancers
The MAnycast2  paper ((Sommese et al., 2020), §5.1) indicated that load balancers might be a cause of FPs when splitting traffic over links. Load balancers typically calculate a hash using packet headers, to determine the outgoing link. To test whether our approach – that varies the ICMP payload and checksum – triggers load balancer decisions, we performed the anycast-based measurement using static probes (i.e., all Workers send exactly the same ICMP Echo Request, without payload and checksum variation). The results match our regular measurement, indicating – as suggested by the literature (Almeida et al., 2020) – that load balancer hashes are calculated only on flow headers. Thus, load balancers do not affect our results as we keep these headers static. This rules out load balancers as a cause of FPs, contradicting the hypothesis in the MAnycast2  paper.
@@ 5.1.5. Benefits of synchronous probing
One shortcoming of MAnycast2  is that it probes an entire hitlist from each VP in sequence, leading to a considerable interval between probes to the same target address. Instead, in MAnycastR we implemented synchronous probing where the Orchestrator ensures Workers send out probes to the same target one second after each other.
To compare our approach to the one used in the MAnycast2 paper (where there was a 13-minute gap between probes to the same target), we performed a measurement using a 13-minute interval between probes. This immediately demonstrates a shortcoming of the MAnycast2 approach: as our deployment has 32 sites, a single run takes around 7 hours. We therefore also ran a second measurement using a 1-minute interval between probes, to obtain a shorter runtime. We compare this to MAnycastR measurements with inter-probe offsets of 1 second and 0 seconds. Figure 4 shows the FP rate for these experiments. It plots the number of FPs on a logarithmic y-axis against the number of VPs receiving on the x-axis. Overall, the number of FPs increases as the inter-probe interval increases. Probing with a 0-second interval yields 13,312 FPs, followed by 14,506 FPs for a 1-second interval, 19,830 FPs for 1 minute, and 198,079 FPs for 13 minutes.
Figure 4. Comparison of MAnycast2  with a 13- and 1- minute interval between probes, and MAnycastR with a 1- and 0-second interval between probes.
An increased probing interval increases the time between first and last probes to a target, leaving more time for route flips to occur. We can therefore likely attribute the rise in FPs for larger intervals to routing instability, where route flips cause targets to respond to a different VP, causing a misclassification as anycast as stated in the MAnycast2 paper (Sommese et al., 2020).
While a reducing the inter-probe interval from 1 to 0 seconds leads to a reduction of 1,194 FPs, we argue the benefits, in terms of responsible probing, when spacing probes (like a regular ping) outweigh the slight decrease in FPs.
@@ 5.1.6. Longitudinal precision
To assess the precision of our daily census we evaluate results for the first 56 days, from March 21 to May 15, 2024. We focus on prefixes detected using ICMPv4 for the anycast-based and GCD approaches. On average, we observe 27.5 k prefixes with the anycast-based approach and 12.1 k prefixes with GCD daily, during this period. Taking the union of all prefixes observed with the anycast-based approach during the 56-day timeframe, we find 78,687 /24-prefixes. Of these prefixes, 15,791 /24s are consistently observed on all 56 days. Conversely 62,896 prefixes only showed up on some days. Of these we find only a few rare cases that are anycast according to GCD as well, but had downtime and/or switched to unicast during the time period. For most we observe they are prefixes not detected as anycast by GCD, and suspect many are FPs. For the GCD-confirmed prefixes we observe 12,605 prefixes detected as anycast for at least one day, 11,359 are observed every day. The 1,246 prefixes not observed every day, include regional anycast deployments that are difficult to detect with GCD, cases of suspected BGP prefix hijacking (causing FPs), and anycast deployments that had downtime.
Overall, we observe that our anycast-based set has a high variability, whereas the GCD set is much more stable, confirming that our combined approach – using anycast and GCD in MAnycastR– is important to achieve high precision.
@@ 5.2. Anycast site enumeration
One feature from iGreedy that we include in MAnycastR is estimating the number of replicas for an anycasted prefix. Figure 5 shows a cumulative distribution plot of the number of anycast sites detected using latency data obtained from Ark (blue) and RIPE Atlas (orange). The circles show the enumeration counts attributed almost entirely to prefixes from three hypergiants (Google, Amazon and Cloudflare).
These counts are a lower bound of the actual number of anycast sites. For example, Google has presence in 103 cities (Google, 2025b) where we find $\sim$41 sites and Cloudflare has PoPs in 300+ cities (Cloudflare, 2025) where we detect $\sim$54 sites. For small deployments (¡ 30 VPs) enumeration is near the actual number of sites which we validate using ground truth in §5.8.
Comparing RIPE Atlas with Ark, we find both have similar results for small anycast deployments (bottom left). For large deployments (top right), RIPE Atlas is able to achieve a higher enumeration of 80 compared to 60 with Ark. However, it has large variability in the number of sites detected due to inconsistency in the number of RIPE Atlas nodes participating in measurements. We provide a more elaborate comparison between these platforms in Appendix A.
@@ 5.3. Increasing coverage
Figure 6. MAnycastR detection of anycast-based method for ICMPv4, TCPv4, and UDPv4.
@@ 5.3.1. Protocol support
To extend coverage compared to iGreedy and MAnycast2 we added TCP and UDP probing. Figure 6 shows a breakdown of which prefixes can be detected using our anycast-based approach with different protocols. First, the bottom left shows the totals for each protocol, i.e., we detect 25,228 anycast prefixes with ICMP, 8,202 with TCP, and 8,192 with UDP. Next, below the histogram we indicate the intersection shown by each bar. We color this by three categories: blue for results of a single protocol, green for groupings of two protocols, and red for the intersection of all three protocols. The bars show the totals and percentage that each intersection gives, e.g., the first bar shows that 12,874 prefixes (48.8% of the total) were discovered to be anycast using only ICMP (i.e., TCP and UDP did not detect these prefixes as anycast). Next, we have the prefixes detected as anycast by both ICMP and UDP, totaling to 4,793, etc.
These results show that ICMP probing uncovers the most anycasted prefixes, as indicated by the ICMP total (bottom-left) and the low totals for the intersections that exclude ICMP (last three bars in the histogram). Furthermore, we see an increase in coverage using TCP and UDP, with 566 prefixes that are only detected as anycast using TCP (third-to-last bar), and 512 with only UDP (second-to-last bar). Of these prefixes detectable only with UDP, we find 97 prefixes received at more than 3 VPs where the anycast-based approach has high accuracy as shown in Table 3. E.g., DNS G-root (operated by the US Department of Defense), LACNIC, Oracle, eBay, several registrars, and various TLDs deploy DNS anycast that is only detectable with anycast-based UDP probing.
Prior work (see §2.3) also attempted to detect and enumerate anycast using CHAOS queries. We therefore added support for CHAOS queries. However, we find CHAOS records are often used for co-located servers at a single location (suggested by values such as ‘auth1’ and ‘auth2’). As such, we find multiple CHAOS records to be a weak indicator of anycast. We provide a more detailed comparison on CHAOS records in Appendix C.
@@ 5.3.2. IPv6 support
To fully cover the IP anycast landscape, we also added support for IPv6. Overall, we find fewer prefixes with the anycast-based approach for IPv6, 6,864 compared to 27,573 for IPv4. This is unsurprising as our IPv6 hitlist is much smaller. Like IPv4, most anycast candidates are detected using ICMP (6,659). Furthermore, we find a large number of anycasted /48s responsive to TCP (4,476). The higher TCP responsiveness, compared to IPv4, was expected due to the different hitlist origins. Unlike the ISI hitlist, which is based on ping-scanning, TUM’s and OpenINTEL’s IPv6 hitlists more likely reflect active services. For the specific distribution for IPv6 we refer to Figure 7 in the appendix.
@@ 5.4. Replicability
Protocol Anycast targets Anycast targets Intersection
@@
Table 4. ICMPv4 anycast-based ATs found using two distinct anycast deployments.
To validate if our methodology is deployment-agnostic, we deploy MAnycastR on an independent, ccTLD registry operated anycast production infrastructure, providing VPs at 12 distinct locations. Using this deployment we perform an anycast-based measurement for ICMPv4 and ICMPv6. The results are shown in Table 4. We observe that the external anycast platform finds fewer candidate anycast prefixes. Especially for IPv4, they find 16,208 IPv4 anycasted /24-prefixes compared to 25,324 prefixes found with our own deployment.
In total, our census finds 11,322 IPv4 anycast targets not found by the ccTLD deployment. In both cases, the vast majority (¿ 98%) are targets that are detected by only 2 VPs. Since the FP rate is high when 2 VPs receive responses, this suggests that the non-intersecting targets are largely FPs.
Of 13,912 IPv4 ATs seen on both platforms, 12,816 are confirmed using GCDArk. Additionally, ATs seen exclusively on our own or on the ccTLD platform include 481 and 112 GCD-confirmed prefixes, respectively. These latter sets of prefixes are small and difficult to detect anycast prefixes only seen by either platform due to topological differences. When taking the union of ATs, we find the anycast-based approach detects 13,409 out of 13,692 GCDArk prefixes (98.0%).
For IPv6 we find 6,996 ATs with our own and 6,501 with the ccTLD platform, with an intersection of 6,255. Most non-intersecting ATs are seen at 2 VPs and not GCD-confirmed.
@@ 5.5.1. Reducing deployment
Measuring the entire hitlist requires 188 M probes for our daily ICMPv4 anycast-based measurement compared to GCDArk requiring 1.3 B probes. For this reason, we perform GCD only towards ATs found with the anycast-based approach. The previous section showed that a smaller anycast deployment finds considerably fewer ATs, thereby reducing the number of targets for the follow-up GCD measurement. However, this comes at a cost of FNs (GCD-confirmed anycast prefixes missed by the anycast-based approach). This raises the question if we can reduce the anycast deployment used in the daily pipeline to reduce the number of ATs with a minimal increase in FNs.
[…]
@@
MAnycastR production 32 VPs 25,324 263 (1.9%) 188 M
GCDArk (full hitlist) 227 VPs 13,692 0 (0.0%) 1,335 M
[…]
To answer this question we performed additional anycast-based measurements. First, we pick two sites per continent, maximizing geographical distance (e.g., sites on the US East and West Coast), totaling 11 VPs (we have one site in Africa). Then, we limit ourselves to the sites that receive the most responses on each continent (6 VPs). Finally, we use only two VPs; one in North America and one in Europe. Table 5 shows the result of these measurements, including the results of our full deployment and the ccTLD deployment.
Interestingly, we observe the highest number of ATs when probing with two sites in each continent. This shows that reducing the number of VPs does not necessarily reduce the number of ATs. Despite the increase in ATs, this deployment has a lower recall due to an increase of 370 FNs compared to our regular anycast-based measurement. When probing with one site per continent, we find 14,221 ATs. However, this comes at a cost of 1,311 missed anycast prefixes. Finally, when probing with only two VPs we find the fewest ATs (12,492) at the cost of the most FNs (2,164). This also shows that even with only two VPs, 84% of GCD-confirmed prefixes are detected, which is unsurprising considering most anycast deployments have a global presence with at least one site in North America and Europe. The FNs of each measurement largely consist of small anycast deployments that are confined to small geographical areas (i.e., regional anycast).

### v3

@@ 5.1.3. Investigating Disagreement
# of sites Candidate $\mathscr{G}$ $\mathscr{M}$ Overlap
receiving anycast (GCD) ($\neg\textbf{GCD}$) (in %)
2 17,996 726 17,270 4.03%
3 597 366 231 61.31%
4 363 309 54 85.12%
5 253 215 38 84.98%
6-10 1,103 954 149 86.49%
11-15 854 713 141 83.49%
16-20 4,608 4,608 0 100.0%
21-25 1,415 1,404 11 99.22%
26-32 3,420 3,420 0 100.0%
Total 30,609 12,715 17,894 41.54%
Table 2. Comparing anycast-based ICMPv4 results per number of sites receiving responses with GCD.
The number of VPs receiving responses for a given target determines whether the anycast-based approach classifies a target as anycast. For a VP count of one, the target is marked as unicast. For a count above one, the target is marked as an anycast candidate. The MAnycast2  paper showed the FP case (unicast responding to multiple sites) most commonly occurs when two sites receive responses. Table 2 compares ‘GCD-confirmed’ and ‘anycast-based only’ (i.e., $\mathscr{G}$ and $\mathscr{M}$ as shown in Figure 3), based on the number of sites receiving responses. We confirm that the vast majority of disagreement occurs when 2 VPs receive responses, followed by fewer cases for 3, 4, and 5 VPs receiving responses.
When 4 or more VPs receive responses they are predominantly confirmed using GCD ($\mathscr{G}$), with only 393 of 12,613 prefixes not confirmed of which 87% originate from Imperva. We suspect this is related to their on-demand DDoS mitigation service, where prefixes are anycast for short periods of time.
Investigating the total prefixes not confirmed by GCD ($\mathscr{M}$), mostly received at 2 and 3 VPs, we observe 12,843 (¿ 70% on any given day) originate from Microsoft’s AS 8075. While we lack information to confirm this, we speculate, based on a 2019 study (Arnold et al., 2019), that this is due to Microsoft’s internal routing policies, where prefixes are announced globally to ingress traffic into their network as soon as possible, with internal routing to a unicast destination. We further suspect that replies egress at multiple edges, based on observed ingress routes, that reach different VPs. Knowing of globally announced prefixes (i.e., global BGP) that route to a single location (i.e., unicast) is valuable as it can help with understanding routing properties for such networks. Using traceroute we confirm probes ingressing at distinct PoPs. Future work is to include global BGP in our census and expand discovery of this practice.
@@ 5.1.4. Influence of load balancers
The MAnycast2  paper ((Sommese et al., 2020), §5.1) indicated that load balancers might be a cause of FPs when splitting traffic over links. Load balancers typically calculate a hash using packet headers, to determine the outgoing link. Literature suggests load balancer hashes are mostly calculated using flow headers (Almeida et al., 2020; Hendriks et al., 2025; Augustin et al., 2007), which we keep static. However, per-packet load balancers may still cause unwanted FPs. To test whether our approach – that varies the ICMP payload and checksum – triggers load balancer decisions, we performed the anycast-based measurement using static probes (i.e., all Workers send exactly the same ICMP Echo Request, without payload and checksum variation). The results match our regular measurement indicating most load balancers do not affect our results as we keep flow headers static.
@@ 5.1.5. Benefits of synchronous probing
The prototype implementation of MAnycast2 probed an entire hitlist from each VP in sequence, leading to 13-minute gaps between probes to the same target. In LACeS we implemented synchronous probing, where Workers send probes to a shared target with one second intervals from each other.
To compare our approach to the one used in the MAnycast2 paper, we performed a measurement using a 13-minute interval between probes. This immediately demonstrates a shortcoming of the MAnycast2 approach: as our deployment has 32 sites, a single run takes around 7 hours. We therefore also ran a second measurement using a 1-minute interval between probes, to obtain a shorter runtime. We compare this to LACeS measurements with inter-probe offsets of 1 second and 0 seconds. Figure 5 shows the the number of prefixes not confirmed using GCD ($\mathscr{M}$), which are mostly FPs, on a logarithmic y-axis against the number of VPs receiving on the x-axis. We show FPs up to 16 VPs receiving as FPs are negligible when more VPs receive replies (see Table 2). Overall, the number of FPs increases as the inter-probe interval increases. Probing with a 0-second interval yields 13,312 FPs, followed by 14,506 FPs for a 1-second interval, 19,830 FPs for 1 minute, and 198,079 FPs for 13 minutes.
Figure 5. FPs found by number of VPs receiving using MAnycast2  with a 13- and 1-minute interval between probes, and LACeS with a 1- and 0-second interval between probes.
An increased probing interval increases the time between first and last probes to a target, leaving more time for route changes to occur. We can therefore likely attribute the rise in FPs for larger intervals to routing instability, causing a misclassification as anycast as stated in the MAnycast2 paper (Sommese et al., 2020). While reducing the inter-probe interval from 1 to 0 seconds leads to a reduction of 1,194 FPs, we argue the benefits, in terms of responsible probing, when spacing probes (like a regular ping) outweigh the slight decrease in FPs.
@@ 5.1.6. Longitudinal precision
We assess the precision of our census by evaluating daily results from census start (March 21, 2024) to date (September 6, 2025) covering 17+ months, focusing on prefixes detected using ICMPv4 for the anycast-based and GCD approaches.
On average, we observed 25.2 k prefixes with the anycast-based approach and 12.9 k prefixes with GCD daily. Taking the union of all prefixes observed with the anycast-based approach during the 534-day timeframe, we find 203 k /24-prefixes. Of these prefixes, 9.9 k (5%) /24s are consistently observed on all 534 days and 193 k (95%) prefixes only showed up on some days. For the latter, we observe they are mostly not detected as anycast by GCD, and suspect many are FPs. For the GCD-confirmed ($\mathscr{G}$) prefixes we observe 16.6 k prefixes detected as anycast for at least one day, 9.7 k (58%) are observed every day. We show in §7 – where we perform a longitudinal analysis – that the combined results of LACeS not only yield a stable census, but also captures both long-term ground-truth confirmed anycast deployments and captures anycast dynamics.
Overall, we observe that our anycast-based set has a high variability, whereas the GCD set is much more stable, confirming that our combined approach – using anycast and GCD in LACeS– is important to achieve high precision.
@@ 5.2. Anycast site enumeration
One feature from iGreedy that we include in LACeS is estimating the number of sites for an anycasted prefix. Figure 6 shows a cumulative distribution plot of the number of sites detected using latency data obtained from Ark (blue) and RIPE Atlas (orange) in Sep. ’24. At the time, Ark measurements were performed using 164 VPs. The circles show the enumeration counts attributed almost entirely to prefixes from Google, Amazon and Cloudflare. These counts are a lower bound of the actual number of anycast sites. For example, Google has presence in 103 cities (Google, 2025b) where we find $\sim$41 sites and Cloudflare has PoPs in 300+ cities (Cloudflare, 2025) where we detect $\sim$54 sites. For small deployments (¡ 30 PoPs) enumeration is near the actual number of sites which we validate using ground truth in §6.
Comparing RIPE Atlas with Ark, we find both have similar results for small anycast deployments (bottom left). For large deployments (top right), RIPE Atlas is able to achieve a higher enumeration of 80 compared to 60 with Ark. However, it has large variability in the number of sites detected due to inconsistency in the number of RIPE Atlas nodes participating in measurements. We provide a more elaborate comparison between these platforms in Appendix B.
@@ 5.3. Increasing coverage
Figure 7. LACeS detection of anycast-based method for ICMPv4, TCPv4, and DNSv4.
@@ 5.3.1. Protocol support
To extend coverage compared to iGreedy and MAnycast2 we added TCP and DNS over UDP probing. Figure 7 shows a breakdown of which prefixes can be detected using our anycast-based approach with different protocols. We detect 30,609 anycast prefixes with ICMP, 11,569 with TCP, and 2,997 with DNS.
The largest group are ICMP-only prefixes, where we find 19,095 prefixes (57.7%) (i.e., TCP and DNS did not detect these prefixes as anycast). Next, we have the prefixes detected as anycast by both ICMP and TCP, totaling to 8,815, etc.
These results show that ICMP probing uncovers the most anycasted prefixes, Furthermore, we see an increase in coverage using TCP and DNS, with 2,161 prefixes that are only detected as anycast using TCP, and 258 with only DNS. For the latter, we find 51 prefixes received at 4 or more VPs where the anycast-based approach has high accuracy as shown in Table 2. E.g., DNS G-root (operated by the US Department of Defense), LACNIC, Oracle, eBay, several registrars, and various TLDs deploy DNS anycast that is only detectable with anycast-based DNS probing. For the 2,161 prefixes only visible with TCP, we confirm 167 with GCD.
Prior work (see §2.3) also attempted to detect and enumerate anycast using CHAOS queries (Fan et al., 2011). We therefore add support for CHAOS queries. However, like prior work we find CHAOS records are often used for co-located servers at a single location (suggested by values such as ‘auth1’ and ‘auth2’) making it a weak indicator of anycast. We provide a more detailed comparison to CHAOS records in Appendix C.
@@ 5.3.2. IPv6 support
Overall, we find fewer prefixes with the anycast-based approach for IPv6, 12,223 compared to 33,068 for IPv4. This is unsurprising as our IPv6 hitlist is much smaller. Like IPv4, most anycast candidates are detected using ICMP (11,840). Furthermore, we find a large number of anycasted /48s responsive to TCP (8,208). The higher TCP responsiveness, compared to IPv4, was expected due to the different hitlist origins. Unlike the ISI hitlist, which is based on ping-scanning, TUM’s and OpenINTEL’s IPv6 hitlists more likely reflect targets running TCP services. For the specific distribution for IPv6 we refer to Figure 14 in the Appendix.
@@ 5.4. Replicability
Protocol Anycast candidates Anycast candidates Intersection
@@
Table 3. ICMP ACs found using two distinct anycast deployments and their intersections.
To validate if our methodology is deployment-agnostic, we deployed LACeS on an independent ccTLD registry-operated anycast production infrastructure providing VPs at 12 distinct locations. Using this deployment we performed an anycast-based measurement in Jan. ’25 for ICMPv4 and ICMPv6. The results are shown in Table 3. We observe that the external anycast platform found fewer candidate anycast prefixes. Especially for IPv4, they found 16,208 IPv4 anycasted /24-prefixes compared to 25,324 prefixes found with our own deployment.
In total, our census found 11,322 IPv4 anycast candidates not found by the ccTLD deployment. For the non-intersecting ACs, the vast majority (¿ 98%) are detected by only 2 VPs. Since the FP rate is high when 2 VPs receive responses, this suggests the non-intersecting candidates are largely FPs.
Of 13,912 IPv4 ACs seen on both platforms, 12,816 (92%) are confirmed using GCD ($\mathscr{G}$). Additionally, ACs seen exclusively on our own or on the ccTLD platform include 481 and 112 GCD-confirmed prefixes, respectively. These latter sets of prefixes are small and difficult to detect anycast prefixes only seen by either platform due to topological differences. In § 5.6 we take a closer look at the impact of routing topology on detection.
For IPv6 we found 6,996 ACs with our own and 6,501 with the ccTLD platform (intersection of 6,255). Most non-intersecting ACs are seen at 2 VPs and not GCD-confirmed.
We also performed anycast-based LACeS measurements using different deployments in May. ’25 First, we deployed LACeS using an additional provider, Melbicom (Melbicom, 2025), that provides 16 VPs from which we announced our IPv4 prefix. This deployment found 18,525 ACs, covering 91.6% of GCD-confirmed prefixes ($\mathscr{G}$) found that day. Manual investigation of several missed GCD-confirmed prefixes shows they are mostly regional to Asia and/or Oceania where Melbicom provides only a single VP. This indicates that LACeS requires multiple VPs in each continent to detect anycast confined within a continent.
To assess whether LACeS can be deployed using multiple hosting providers we announced our IPv4 prefix using both Vultr (as used for our daily census) and Melbicom leading to 48 VPs combined. In total, we find 31,340 ACs covering 95.7% of GCD-confirmed prefixes ($\mathscr{G}$) showing operators can deploy LACeS on diverse anycast deployments. Surprisingly, we find more GCD-confirmed prefixes with the Vultr only deployment (96.2%).
These results show that LACeS is deployment-agnostic and can also be deployed on anycast infrastructures that make use of multiple hosting providers. We encourage operators to run their own LACeS census. Additionally, we find detection of regional anycast is sensitive to topological connectivity of the deployment used.
A more detailed investigation of the effects of topology (e.g., upstream providers per PoP) may yield a better understanding as to when and why the FN and FP cases of the anycast-based approach occur. In particular, we can use, e.g., reverse traceroute (Katz-Bassett et al., 2010) to find multi-path topologies that cause unicast targets responses to reach multiple sites.
@@ 5.5.1. Reducing deployment
Measuring the hitlist requires 188 M probes for our daily ICMPv4 anycast-based measurement compared to the GCDLS from Dec. ’24 requiring 1.3 B probes. For this reason, we only perform GCD towards ACs found with the MAnycast2 approach. The previous section showed that a smaller anycast deployment finds considerably fewer ACs, thereby reducing the number of targets for the follow-up GCD measurement. However, this comes at a cost of FNs (GCD-confirmed anycast prefixes missed by the anycast-based approach). This raises the question if we can reduce the anycast deployment used in the daily pipeline to reduce the number of ACs with a minimal increase in FNs.
Deployment ACs $\neg\textbf{GCD\textsubscript{{LS}}}$ ($\neg\textbf{GCD\textsubscript{{LS}}}$ %) Cost
@@
Melbicom 16 VPs 18,525 1,155 (8.4%) 94 M
TANGLED (Bertholdo et al., 2021) (Vultr) 32 VPs 25,396 524 (3.8%) 189 M
Vultr+Melbicom 48 VPs 31,340 593 (4.3%) 283 M
GCDLS Dec. ’24 (full) 227 VPs 13,692 0 (0.0%) 1,335 M
Table 4. The number of anycast candidates (ACs), missed GCDLS prefixes ($\neg\textbf{GCD\textsubscript{{LS}}}$) and probing cost for various anycast deployments and GCDLS from Dec. ’24 (full hitlist).
To answer this question we performed additional anycast-based measurements. First, we pick two sites per continent, maximizing geographical distance (e.g., sites on the US East and West Coast), totaling 11 VPs (we have one site in Africa). Then, we limit ourselves to the sites that receive the most responses on each continent (6 VPs). Finally, we use only two VPs; one in North America and one in Europe. Table 4 shows the result of these measurements, including the results of the ccTLD, Vultr (as used for TANGLED), Melbicom, and Vultr + Melbicom deployments.
Interestingly, we observe the second largest number of ACs when probing with two sites in each continent. This shows that reducing the number of VPs does not necessarily reduce the number of ACs. Despite the increase in ACs, this deployment has a lower recall due to an increase of FNs compared to TANGLED. When probing with one site per continent, we find 14,221 ACs. However, this comes at a cost of 1,311 missed anycast prefixes. Finally, when probing with only two VPs we find the fewest ACs (12,492) at the cost of the most FNs (2,164). This also shows that even with only two VPs, 84% of GCDLS prefixes are detected, which is unsurprising considering most anycast deployments have a global presence with at least one site in North America and Europe as we show in §6. The FNs of each measurement largely consist of small anycast deployments that are confined to small geographical areas (i.e., regional anycast).